# A multi-task deep learning approach for lane-level pavement performance prediction with segment-level data

### v1

@@ 5 CONCLUSION
Lane-level performance predictions offer more granular information for developing subsequent maintenance strategies and creating maintenance plans compared to segment-level predictions. The paper proposed a multi-task deep learning approach, accounting for the similarities and differences in pavement performance time series and auxiliary features of different lanes, for lane-level pavement performance prediction based on segment-level pavement inspection data. It has undergone rigorous validation and ablation experiments using segment-level pavement inspection data from Zhengzhou, Jiaozuo, and Luoyang in Henan Province, China. The principal conclusions are presented below.
(1) Our model performs better in predicting PCI, PQI, and RQI based on three distinct datasets. It illustrates the efficacy of multitask-learning based model in leveraging related tasks to improve overall and task prediction precision in pavement performance prediction.
(2) Increasing the number of lanes predicted simultaneously can improve the prediction accuracy of the overall and each task and achieve higher performance improvement compared to other models.
In addition, there are still two issues deserving in-depth discussions. First, the spatial relationship of the prediction units should be further explored during the prediction. Since the prediction units of adjacent positions usually have similar and correlated performance. Second, very few prediction units may happen sharp drops in pavement performance, which results in worse prediction accuracy.

### v2

@@ 5 CONCLUSION
The lane-level pavement performance prediction offers high-resolution information for developing maintenance strategies and plans. However, there is a lack of high-resolution lane-level pavement performance data generally. Thus, this study proposes a multi-task deep learning approach that can predict pavement performance of multiple lanes in one unified model with enough historical segment-level data. The proposed deep learning approach mixed LSTM and ANN with a multi-task learning structure. Such design allows us to capture temporal evolution in pavement performance at both segment-level and lane-level, but also unify multiple prediction tasks into one modeling structure. Rigorous validation and ablation experiments have been conducted utilizing real pavement performance inspection data obtained from three cities within Henan Province, China: Zhengzhou, Jiaozuo, and Luoyang.
(1) The proposed model achieves an average of 10.309%, 5.816%, and 4.936% MAPE when predicting PCI, PQI, and RQI on the 2-lane, 3-lane, and 4-lane datasets and gets an average absolute maximum reduction in MAPE of 2.774%, 2.699%, and 3.092%, compared with the ’lane-specific model’ and ’mix model’. Besides, the model gets the lowest MAPE of each lane in most cases.
(2) Ablation experiments on model structure show the LSTM-based shared layer, multiple task-specific heads, and horizontal concatenation process play an important supporting role in reducing prediction error, and the reduction are 1.631%, 1.611%, and 1.806% when predicting PCI, PQI, and RQI, respectively. Ablation experiments on pavement auxiliary features show all the utilized auxiliary features contribute positively to enhancing the accuracy of the predictions. However, there is not one feature that will greatly improve prediction accuracy. The max reduction of MAPE is 0.547%, 0.697%, and 0.663% when predicting PCI, PQI, and RQI.
(3) In most cases, the MAPE of each lane is the lowest using the proposed model compared with the lane-specific and mix models. The outcomes show the effectiveness of the proposed model in leveraging related tasks to enhance the overall and lane-specific forecasting accuracy in pavement performance analysis. As the number of lanes increases, the difference in MAPE among lanes narrows. However, it is worth noting that increasing the number of tasks does not necessarily increase the magnitude of MAPE reduction.
Future research may encompass in-depth discussions on two unresolved issues. First, the spatial relationship of the prediction units should be further explored during the prediction. Since the adjacent prediction units, or two units with similar environment, usually have similar pavement deterioration patterns. Second, few road units may experience very sharp drops in pavement performance, which violates most common deterioration patterns. This is challenging in precisely predicting such sharp drops with almost all models in the literature.
Author contributions
Bo Wang: Writing – original draft, Visualization, Validation, Software, Methodology, Investigation. Wenbo Zhang: Writing – review & editing, Conceptualization, Methodology, Supervision, Resources, Project administration. Yunpeng Li: Methodology.
Funding sources
This work was supported by the National Key Research and Development Program of China [grant number 2022YFB4300300]